1) Give yourself 100# for the 5 feet plus 5# for every inch over 5 feet tall. Using this description. What height corresponds to a recommended weight of 135#?
Answer by ewatrrr(24785)


Hi
5# for every inch over 5 feet
Let x represent the number of inches over 5 ft
135lbs = 100 + 35lb = 100 + 5x
35 = 5x
7 = x, the number of inches over 5 ft
Height would be 5'7" for a recommended weight of 135#?
CHECKING our Answer***
5'7" translates to 100lb + 7*5lb OR 135lb of recommended weight